### v2

# How to build transfer matrices, one wave at a time

Abstract
We show how to build the closed-form expression of transfer matrices for wave propagation in layered media. The key is to represent the propagation across the piece-wise constant medium as a superposition of a finite number of paths ($2^{N-1}$ paths for a medium with $N$ layers), each one of them contributing a certain phase change (corresponding to signed sums of the phase change in each individual layer) and amplitude change (corresponding to the pattern of transmission and/or reflection associated to each path). The outlined technique is combinatorial in nature: it begins with the linear governing equations in frequency domain, whose fundamental solution is known, then it enumerates the finite number of paths across the overall system, then computes their associated phase and amplitude change, and finally adds all the possible paths to find the final result. Beyond providing physical insight, this “path-by-path” construction can also circumvent the need for transfer matrix numerical multiplication in many practical applications, potentially enabling substantial computational savings.
1 Intro
Understanding wave propagation in layered systems is essential for designing and optimizing a wide range of subjects, such as phononic [1] and photonic crystals [2], elastodynamics [3], acoustic metamaterials [4], and quantum well lattices [5]. These systems often rely on engineered interference, bandgap formation, or resonance phenomena that emerge from the interaction of superimposing waves. Logically, accurately predicting how waves reflect, transmit, or localize within such architectures is key to tailoring functional properties — whether it is achieving sound insulation [6, 7], controlling light flow [8] or design of seismic barriers [9]. Performing these predictive calculations, especially in complex or highly heterogeneous systems, relies on the transfer matrix approach [10, 11, 12, 13, 14], which is both mathematically and physically sound, and robust and efficient from the computational implementation standpoint.
Both “transfer” matrices and “scattering” matrices are used to analyze wave propagation in layered media across fields like photonics, phononics, acoustics, elastodynamics, geotechnical earthquake engineering and seismology (the term “propagator” or “propagation matrix” is also used [15, 16]). The difference between the two matrices is that while transfer matrices relate wave amplitudes across interfaces layer by layer (making them ideal for modeling periodic structures and computing dispersion relations), scattering matrices relate incoming to outgoing wave amplitudes at a boundary or structure (thus providing direct access to reflection and transmission coefficients). Transfer matrices are widely used for forward modeling and band structure analysis, whereas scattering matrices are employed in inverse problems.
To our knowledge, there have not been many attempts to derive closed-form expressions (let alone to provide a physical interpretation) for the general form of the scattering matrix, the transfer matrices, or of the latter’s trace. Shen and Cao presented the general expression of the trace of $2\times 2$ transfer matrix entries using a finite product of cosines times a finite sum of cosines and sines [17] in the context of acoustic dispersion relations; in an appendix, they showed the explicit expression for two, four and six layers. Vinh and colleagues [18] have presented the expression of the four entries of the $2\times 2$ matrices that appear in acoustics, optics and elastodynamics, aiming to provide a general expression of the reflection and transmission coefficient across the full stack of $N$ isotropic homogeneous layers. No mention to “harmonics” was made before Ref.[19], which introduced the “harmonic decomposition of the trace”, a result that has already been exploited for both phononics [20] and geotechnical earthquake engineering [21]. We will show here that this harmonic decomposition is actually a “path-wise decomposition”. This is the first attempt to infuse physical meaning to all the aforementioned expressions obtained in the course of analyzing wave propagation in layered media.
2 General matrix expressions
Staring from the PDE that governs wave propagation, through either Fourier transform or a plane-wave expansion and dividing the problem in frequency domain into “amplitude” and “gradient”, one reached a linear first-order vector ODE. Consider a perturbation propagating across a homogeneous medium in the $x$-direction, let $\boldsymbol{\mathsf{f}}(x,\omega)=[f(x,\omega)\,,f^{\prime}(x,\omega)]^{\top}$, where $f(\omega)$ is the amplitude of the field of interest and $f^{\prime}(\omega)$ its gradient. Its propagation along $x$ is governed by the map
$\begin{bmatrix}f\\ f^{\prime}\end{bmatrix}_{x=x_{i+1}}=\boldsymbol{T}(x_{i+1},x_{i})\begin{bmatrix}f\\ f^{\prime}\end{bmatrix}_{x=x_{i}}=\begin{bmatrix}\cos\left({\mathrm{k}_{i}l_{i}}\right)&{\sin\left({\mathrm{k}_{i}l_{i}}\right)\over\mathrm{k}_{i}}\\ -{\mathrm{k}_{i}\sin\left({\mathrm{k}_{i}l_{i}}\right)}&\cos\left({\mathrm{k}_{i}l_{i}}\right)\end{bmatrix}\begin{bmatrix}f\\ f^{\prime}\end{bmatrix}_{x=x_{i}}\,,$ (1)
where $l_{i}=x_{i+1}-x_{i}$ is the thickness of the $i$-th homogeneous layer and $\mathrm{k}_{i}$ is the wavenumber in it. To understand situations where eq. 1 appear, see Table 1. To propagate a state across the piecewise-defined medium, just chain successive matrix multiplications:
$\begin{bmatrix}f\\ f^{\prime}\end{bmatrix}_{x=x_{N}}=\underbrace{\boldsymbol{T}_{N}(x_{N},x_{N-1})\ldots\boldsymbol{T}_{k}(x_{k},x_{k-1})\ldots\boldsymbol{T}_{1}(x_{1},x_{0})}_{\boldsymbol{T}}\begin{bmatrix}f\\ f^{\prime}\end{bmatrix}_{x=x_{0}}\,.$ (2)
$\boldsymbol{T}$ is called “cumulative” transfer matrix, and it is the product of $N$ “atomic” transfer matrices as in eq. 1, each one associated to a homogeneous piece of the heterogeneous medium.
What does this series of matrix multiplication mean physically? Understanding of the structure of the cumulative transfer matrix is critical, for a myriad reason: its trace governs the dispersion relation in 1D both phononics [22] and photonics [8], some of its entries can be related to transfer matrices [15] and HV ratios [23] in seismology and geotechnical earthquake engineering, or it can be related to the evolution of wavefunctions [24]. The scattering matrix appears, e.g., in seismology [16] or other disciplines whenever we want to compute the relation between incoming, reflecting and wave amplitudes.
While the transfer matrix formulation as in eq. 1 works on amplitudes at each interface, the scattering matrix looks at amplitudes propagating in and out of the stack. To find the relation between the two, one can write: $f=Ae^{-\mathrm{i}\mathrm{k}x}+Be^{\mathrm{i}\mathrm{k}x}$, where $A,\,B$ are the amplitudes of the right-propagating wave and the left-propagating one, respectively, and $\mathrm{i}$ is the imaginary unit. In the i-th layer:
$\begin{bmatrix}A\\ B\end{bmatrix}_{x=x_{i}}=\begin{bmatrix}e^{\mathrm{i}kx_{i}}/2&\mathrm{i}e^{\mathrm{i}\mathrm{k}x_{i}}/2k\\ e^{-\mathrm{i}\mathrm{k}x_{i}}/2&-\mathrm{i}e^{-\mathrm{i}\mathrm{k}x_{i}}/2k\end{bmatrix}\begin{bmatrix}f\\ f^{\prime}\end{bmatrix}_{x=x_{i}}=\boldsymbol{E}(x_{i})\begin{bmatrix}f\\ f^{\prime}\end{bmatrix}_{x=x_{i}}\,,$ (3)
where $\boldsymbol{E}$ is basis matrix for a second-order linear wave equation (or Helmholtz-type problem) in 1D. hence, the amplitudes of the outgoing waves (at $x=x_{n}$) in terms of the incoming waves (at $x=x_{0}$):
$\begin{bmatrix}A\\ B\end{bmatrix}_{x=x_{N}}=\boldsymbol{E}(x_{N})\boldsymbol{T}(x_{N},\,x_{0})\boldsymbol{E}(x_{0})^{-1}\begin{bmatrix}A\\ B\end{bmatrix}_{x=x_{0}}=\boldsymbol{S}(x_{N},\,x_{0})\begin{bmatrix}A\\ B\end{bmatrix}_{x=x_{0}}\,,$ (4)
where $\boldsymbol{S}$ is the scattering matrix. See how, once the transfer matrix of the medium is ready, the scattering matrix is just two simple matrix multiplications away.
System Wavenumber (k) Impedance (Z)
1D Shear Waves $\dfrac{\omega}{\sqrt{G/\rho}}$ $\sqrt{G\rho}$
1D Pressure Waves $\dfrac{\omega}{\sqrt{B/\rho}}$ $\sqrt{\rho B}$
Electromagnetic Waves $\omega\sqrt{\mu\epsilon}$ $\sqrt{\mu\over\epsilon}$
Table 1: Wavenumber expressions for various 1D wave systems. $\omega$ represents the frequency all the wave in all cases. $G$ is the shear modulus of an elastic medium and $B$ is the bulk modulus of an acoustic one. $\mu$ is the permeability and $\epsilon$ the permittivity of the medium.
3 Building the transfer matrix
3.1 Physical interpretation of propagation in one layer
[…]
$T_{11}=\cos\left({\mathrm{k}}x\right)={1\over 2}e^{\mathrm{i}{\mathrm{k}x}}+{1\over 2}e^{-\mathrm{i}{\mathrm{k}x}}={1\over 2}e^{\mathrm{i}{\omega x\over c}}+{1\over 2}e^{-\mathrm{i}{\omega x\over c}}\,,$ (5)
where $c$ is the signal propagation velocity and each exponential corresponds to a phase shift ±$\pm\omega x/c$, i.e., the wave oscillating, each harmonic with its own frequency $\omega$, and moving from 00 to $x$. This is all too familiar to anyone who has ever tried to propagate an initial 1D wavelet: at $t>0$, the wave package divides its original amplitude in two halves, and each portion propagates in opposite directions but at the same velocity for all the harmonics whose superposition render the waveform.
So $T_{11}$ would give us the new amplitudes of $f$ created by an impulse, what about $T_{21}$? From the transfer matrix, it follows that it must represent the wave gradient associated to those two new waves in eq. 5:
$T_{21}={\mathrm{i}\mathrm{k}\over 2}e^{\mathrm{i}{\mathrm{k}x}}-{\mathrm{i}\mathrm{k}\over 2}e^{-\mathrm{i}{\mathrm{k}x}}=-\mathrm{k}\sin\left({\mathrm{k}x}\right)\,,$ (6)
as expected, what confirms the interpretation as the gradient created by the impulse in $\boldsymbol{\mathsf{f}}=[1\,,0]^{\top}$. Exactly the same procedure can be used for an impulse in the gradient $\boldsymbol{\mathsf{f}}=[0\,,1]^{\top}$, what would render an interpretation for $T_{12}$ as the amplitude created by this load and $T_{22}$ as its gradient.
Once this exercise has revealed the physical meaning of the transfer matrix for one constant piece of medium, nothing deters us from employing the same arguments for more layers.
3.2 Two-layer transfer matrix
In this case, $\boldsymbol{T}$ must be able to represent the state at the edge of the second layer. The first layer is indexed with “1” and the second one with “2”. Consider the impulse originating on the edge of layer 1, it would propagate perpendicularly to the interface until reaching it, there we know that it will be partially transmitted and partially reflected, according to coefficients $\mathsf{T}_{1\to 2}$ and $\mathsf{R}$ satisfying $1+\mathsf{R}_{1\to 2}=\mathsf{T}_{1\to 2}$. Since we are impinging from 1 to 2, the transmission/reflection coefficients are
$\mathsf{T}_{1\to 2}={2Z_{2}\over Z_{1}+Z_{2}}={2\over\left(1+{Z_{1}\over Z_{2}}\right)}\,,\quad\mathsf{R}_{1\to 2}={Z_{1}-Z_{2}\over Z_{1}+Z_{2}}={\mathsf{T}_{1\to 2}\over 2}\left(1-{Z_{1}\over Z_{2}}\right)\,,$ (7)
where $Z_{i}$ is the corresponding impedance of each medium (see Table 1).
Each portion of wave originating at the left edge of 1 will accumulate phase proportional to the trip across 1, then it will interact with both reflection and transmission associated with 2, adding or subtracting phase.
Let us compute $T_{11}$. Introducing $t_{1}=l_{1}/c_{1}$ the time spent to propagate across layer 1 and $t_{2}=l_{2}/c_{2}$ the time spent to propagate across layer 2, we define two possible paths:
  • 1.
    Path I: let us call this path “$[+1,+1]$”, meaning “propagation in 1, transmission in 2”. The phase change would be $\omega(t_{1}+t_{2})$, proportion of total amplitude would be $1/\mathsf{T}_{1\to 2}$. It would correspond to
    ${1\over\mathsf{T}_{1\to 2}}\left({1\over 2}e^{\mathrm{i}\omega(t_{1}+t_{2})}+{1\over 2}e^{-\mathrm{i}\omega(t_{1}+t_{2})}\right)={1\over\mathsf{T}_{1\to 2}}\cos\left(\omega(t_{1}+t_{2})\right)$ (8)
  • 2.
    Path II: let us call this path “$[+1,-1]$”, meaning “propagation in 1, reflection in 2”. The phase change would be $\omega(t_{1}-t_{2})$, the proportion of total amplitude would be $\mathsf{R}_{1\to 2}/\mathsf{T}_{1\to 2}$. It would correspond to
    ${\mathsf{R}_{1\to 2}\over\mathsf{T}_{1\to 2}}\left({1\over 2}e^{\mathrm{i}\omega(t_{1}-t_{2})}+{1\over 2}e^{-\mathrm{i}\omega(t_{1}-t_{2})}\right)={\mathsf{R}_{1\to 2}\over\mathsf{T}_{1\to 2}}\cos\left(\omega(t_{1}-t_{2})\right)$ (9)
Thus, $T_{11}$, amplitude response to an amplitude impulse at the initial edge of the medium, summing the two possible interference paths:
$T_{11}={1\over\mathsf{T}_{1\to 2}}\cos\left(\omega(t_{1}+t_{2})\right)+{\mathsf{R}_{1\to 2}\over\mathsf{T}_{1\to 2}}\cos\left(\omega(t_{1}-t_{2})\right)\,.$ (10)
Expanding the sum in the cosines’ argument and substituting the coefficients in terms of the impedances eq. 7 in Equation 10, we reach the form that is usually found in the literature:
$T_{11}$ $=\cos\left(\omega t_{1}\right)\cos\left(\omega t_{2}\right)-{Z_{1}\over Z_{2}}\sin\left(\omega t_{1}\right)\sin\left(\omega t_{2}\right)\,.$ (11)
For $T_{22}$ we can use the same arguments (gradient response to a gradient impulse at the initial edge of the medium), just recall to use the proper transmission/reflection coefficients to relate to gradient amplitudes:
$\mathsf{T}_{1\to 2}^{\prime}={Z_{1}\over Z_{2}}\mathsf{T}_{1\to 2}={2\over\left(1+{Z_{2}\over Z_{1}}\right)}\,,\quad\mathsf{R}_{1\to 2}^{\prime}=-\mathsf{R}_{1\to 2}={\mathsf{T}_{1\to 2}^{\prime}\over 2}\left({1-{Z_{2}\over Z_{1}}}\right)\,.$ (12)
Hence,
$T_{22}$ $={1\over\mathsf{T}_{1\to 2}^{\prime}}\cos\left(\omega(t_{1}+t_{2})\right)+{\mathsf{R}_{1\to 2}^{\prime}\over\mathsf{T}_{1\to 2}^{\prime}}\cos\left(\omega(t_{1}-t_{2})\right)$ (13a)
$=\left(1+{Z_{2}\over Z_{1}}\right)\cos\left(\omega(t_{1}+t_{2})\right)+\left(1-{Z_{2}\over Z_{1}}\right)\cos\left(\omega(t_{1}-t_{2})\right)$ (13b)
$=\cos\left(\omega t_{1}\right)\cos\left(\omega t_{2}\right)-{Z_{2}\over Z_{1}}\sin\left(\omega t_{1}\right)\sin\left(\omega t_{2}\right)\,.$ (13c)
Likewise, $T_{21}$ represents the gradient associated to an amplitude impulse at the end of the second layer:
$T_{21}$ $=-\mathrm{k}_{2}{1\over\mathsf{T}_{1\to 2}}\sin\left(\omega(t_{1}+t_{2})\right)-\mathrm{k}_{2}{\mathsf{R}_{1\to 2}\over\mathsf{T}_{1\to 2}}\sin\left(\omega(t_{1}-t_{2})\right)$ (14a)
$=-\mathrm{k}_{2}\sin(\mathrm{k}_{2}l_{2})\cos(\mathrm{k}_{1}l_{1})-\mathrm{k}_{1}\sin(\mathrm{k}_{1}l_{1})\cos(\mathrm{k}_{2}l_{2})\,.$ (14b)
The second equivalent form is the one in which it appears if we do the transfer matrix multiplication. The other non-diagonal term corresponds to the amplitude in the last layer associated to the gradient impulse:
$T_{12}$ $=-{1\over\mathrm{k}_{2}}{1\over\mathsf{T}_{1\to 2}^{\prime}}\sin\left(\omega(t_{1}+t_{2})\right)-{1\over\mathrm{k}_{2}}{\mathsf{R}_{1\to 2}^{\prime}\over\mathsf{T}_{1\to 2}^{\prime}}\sin\left(\omega(t_{1}-t_{2})\right)$ (15)
$={\sin(\mathrm{k}_{2}l_{2})\over\mathrm{k}_{2}}\cos(\mathrm{k}_{1}l_{1})+{\sin(\mathrm{k}_{1}l_{1})\over\mathrm{k}_{1}}\cos(\mathrm{k}_{2}l_{2})\,.$ (16)
Thus we have built the four entries of the transfer matrix for two layers (10, 13a, 14a, 14a); had we done the elementary matrix multiplication we would have got directly (11, 13c, 14b, 16).
3.3 Three-layer transfer matrix
For further illustration, let us compute $T_{11}$ by counting and then combining all the possible paths in the 3-layer case:
  • 1.
    Path I $[+1,+1,+1]$: “propagation, transmission, transmission”, phase change $\omega(t_{1}+t_{2}+t_{2})$, proportion of total amplitude $1/\mathsf{T}_{1\to 2}\times 1/\mathsf{T}_{2\to 3}$. It would correspond to
    ${1\over\mathsf{T}_{1\to 2}}{1\over\mathsf{T}_{2\to 3}}\cos\left(\omega(t_{1}+t_{2}+t_{3})\right)$ (17)
  • 2.
    Path II $[+1,+1,-1]$: “propagation, transmission, reflection”, phase change $\omega(t_{1}+t_{2}-t_{3})$, proportion of total amplitude $1/\mathsf{T}_{1\to 2}\times\mathsf{R}_{2\to 3}/\mathsf{T}_{2\to 3}$. It would correspond to
    ${1\over\mathsf{T}_{1\to 2}}{\mathsf{R}_{2\to 3}\over\mathsf{T}_{2\to 3}}\cos\left(\omega(t_{1}+t_{2}-t_{3})\right)$ (18)
  • 3.
    Path III $[+1,-1,+1]$: “propagation, reflection, transmission”, phase change $\omega(t_{1}-t_{2}+t_{3})$, proportion of total amplitude $\mathsf{R}_{1\to 2}/\mathsf{T}_{1\to 2}\times 1/\mathsf{T}_{2\to 3}$. It would correspond to
    ${\mathsf{R}_{1\to 2}\over\mathsf{T}_{1\to 2}}{1\over\mathsf{T}_{2\to 3}}\cos\left(\omega(t_{1}-t_{2}+t_{3})\right)$ (19)
  • 4.
    Path IV $[+1,-1,-1]$: “propagation, reflection, reflection”, phase change $\omega(t_{1}-t_{2}-t_{3})$, proportion of total amplitude $\mathsf{R}_{1\to 2}/\mathsf{T}_{1\to 2}\times\mathsf{R}_{2\to 3}/\mathsf{T}_{2\to 3}$. It would correspond to
    ${\mathsf{R}_{1\to 2}\over\mathsf{T}_{1\to 2}}{\mathsf{R}_{2\to 3}\over\mathsf{T}_{2\to 3}}\cos\left(\omega(t_{1}-t_{2}-t_{3})\right)\,.$ (20)
These are all the four possible paths in a three-piece system. Adding all up:
$T_{11}$ $={1\over\mathsf{T}_{1\to 2}}{1\over\mathsf{T}_{2\to 3}}\cos\left(\omega(t_{1}+t_{2}+t_{3})\right)+{1\over\mathsf{T}_{1\to 2}}{\mathsf{R}_{2\to 3}\over\mathsf{T}_{2\to 3}}\cos\left(\omega(t_{1}+t_{2}-t_{3})\right)$
$+{\mathsf{R}_{1\to 2}\over\mathsf{T}_{1\to 2}}{1\over\mathsf{T}_{2\to 3}}\cos\left(\omega(t_{1}-t_{2}+t_{3})\right)+{\mathsf{R}_{1\to 2}\over\mathsf{T}_{1\to 2}}{\mathsf{R}_{2\to 3}\over\mathsf{T}_{2\to 3}}\cos\left(\omega(t_{1}-t_{2}-t_{3})\right)\,.$ (21)
It is illustrative to remark that all the coefficients of the cosines add up to 1:
${1\over\mathsf{T}_{1\to 2}}{1\over\mathsf{T}_{2\to 3}}+{1\over\mathsf{T}_{1\to 2}}{\mathsf{R}_{2\to 3}\over\mathsf{T}_{2\to 3}}+{\mathsf{R}_{1\to 2}\over\mathsf{T}_{1\to 2}}{1\over\mathsf{T}_{2\to 3}}+{\mathsf{R}_{1\to 2}\over\mathsf{T}_{1\to 2}}{\mathsf{R}_{2\to 3}\over\mathsf{T}_{2\to 3}}$
$={1\over\mathsf{T}_{1\to 2}}\left(\left[{1\over\mathsf{T}_{2\to 3}}+{\mathsf{R}_{2\to 3}\over\mathsf{T}_{2\to 3}}\right]+\mathsf{R}_{1\to 2}\left[{1\over\mathsf{T}_{2\to 3}}+{\mathsf{R}_{2\to 3}\over\mathsf{T}_{2\to 3}}\right]\right)$
$={1\over\mathsf{T}_{1\to 2}}\left(1+\mathsf{R}_{1\to 2}\right)=1\,.$
Compare the result to the one reported in the literature (obtained from direct elementary transfer matrix multiplication):
$T_{11}=\cos(\omega t_{1})\cos(\omega t_{2})$ $\cos(\omega t_{3})\Bigg{(}1-{Z_{1}\over Z_{2}}\tan(\omega t_{1})\tan(\omega t_{2})$
$+{Z_{1}\over Z_{3}}\tan(\omega t_{1})\tan(\omega t_{3})-{Z_{2}\over Z_{3}}\tan(\omega t_{2})\tan(\omega t_{3})\Bigg{)}\,.$ (22)
Expand the product to eliminate the tangent terms, use the product-to-sum trigonometric identities to turn the products of sines and cosines into cosines, plus recognize the coefficients in terms of $\mathsf{T}_{1\to 2},\mathsf{R}_{1\to 2}$ and $\mathsf{T}_{2\to 3},\mathsf{R}_{2\to 3}$ to recover eq. 21. The rest of the entries can be obtained in like fashion.
3.4 N-layer transfer matrix
As hinted by the previous examples, each new layer adds a new interface, which doubles the number possible paths, therefore for $N$ layers and $N-1$ interfaces the number of paths to account for is $2^{N-1}$. The $j$-th path can be labeled symbolically with a vector $\boldsymbol{\mathrm{e}}_{\text{j}}$ containing $N$ entries (one per layer), “+1” if the initial wave is partially transmitted or “-1” reflected at each interface, and which can be seen the j-th row of a $N\times 2^{N-1}$ matrix $\boldsymbol{\mathrm{e}}$. So, we can always write $T_{11}$ as a sum over all paths and obtain:
$T_{11}=\sum_{{j}=1}^{2^{N-1}}\mathcal{T}_{j}\cos(\tau_{j}\omega)\ ,$ (23)
where
$\tau_{j}\omega=\omega\boldsymbol{\mathrm{e}}_{j}\cdot\boldsymbol{t}$ (24)
is the total phase change across each path ($\boldsymbol{t}$ is the vector of times in each layer, i.e., $l_{i}/c_{i}$ for $i=1,\ldots,N$), and the total amplitude change is therefore
$\mathcal{T}_{j}=2\prod_{i=1}^{N}{1\over 2}\left(1+[\boldsymbol{\mathrm{e}}_{j}]_{i}{Z_{i}\over Z_{i+1}}\right)\,.$ (25)
See how the combinatorial structure of the problem is encoded in a single matrix $\boldsymbol{\mathrm{e}}$, which is reminiscent of Hadamard matrices and of truth tables.
For $T_{22}$, one gets the same as eq. 23, the only difference (arising from the use of different transmission/reflection coefficients for the gradient) is that in the path coefficients in this case, call them $\mathcal{T}_{j}^{\prime}$ are obtained from replacing $Z_{i}/Z_{i+1}$ by $Z_{i+1}/Z_{i}$ in eq. 25. This also means that whenever we want to build the trace of $\boldsymbol{T}$, we can also use eq. 23, replacing $Z_{i}/Z_{i+1}$ in eq. 25 by $(Z_{i}/Z_{i+1}+Z_{i+1}/Z_{i})$ [19, 20].
The anti-diagonal terms $T_{21}$ and $T_{12}$ can be obtained, respectively, from localizing the derivative of $T_{11}$ at the end of the last layer, and localizing the integral of $T_{22}$ at the last layer:
$T_{21}$ $=-\mathrm{k}_{N}\sum_{{j}=1}^{2^{N-1}}\mathcal{T}_{j}\sin(\tau_{j}\omega)\ ,$ (26a)
$T_{12}$ $={1\over\mathrm{k}_{N}}\sum_{{j}=1}^{2^{N-1}}\mathcal{T}_{j}^{\prime}\sin(\tau_{j}\omega)\ ,$ (26b)
Even before presenting here the physical logic that allows building each entry of the transfer matrix path by path, we knew the expressions (23, 25, 24, 26a, 26b) to hold, as they have already been derived in a purely mathematical tour de force: the first general form of the trace of $\boldsymbol{T}$ was derived exploiting the equivalence between transfer matrices and Möbius transformations [19]. The non-diagonal terms were presented in Ref. [21], and an alternative proof, relying solely on tracking transfer matrix products, has been published in Ref. [20].
3.5 N-layer scattering matrix
Having the general form of all the entries of the $N$-layer transfer matrix, we can also build the $N$-layer scattering matrix using eq. 4:
$(S_{11})2\mathrm{k}_{N}e^{-\mathrm{i}(\mathrm{k}_{N}l_{N}-\mathrm{k}_{1}l_{1})}$ $=\mathrm{k}_{N}T_{11}+\mathrm{k}_{1}T_{22}+\mathrm{i}\left(T_{21}-\mathrm{k}_{1}\mathrm{k}_{N}T_{12}\right)\,,$ (27a)
$(S_{22})2\mathrm{k}_{N}e^{\mathrm{i}(\mathrm{k}_{N}l_{N}+\mathrm{k}_{1}l_{1})}$ $=\mathrm{k}_{N}T_{11}+\mathrm{k}_{1}T_{22}-\mathrm{i}\left(T_{21}-\mathrm{k}_{1}\mathrm{k}_{N}T_{12}\right)\,,$ (27b)
$(S_{12})2\mathrm{k}_{N}e^{-\mathrm{i}(\mathrm{k}_{N}l_{N}-\mathrm{k}_{1}l_{1})}$ $=\mathrm{k}_{N}T_{11}-\mathrm{k}_{1}T_{22}+\mathrm{i}\left(T_{21}+\mathrm{k}_{1}\mathrm{k}_{N}T_{12}\right)\,,$ (27c)
$(S_{21})2\mathrm{k}_{N}e^{\mathrm{i}(\mathrm{k}_{N}l_{N}+\mathrm{k}_{1}l_{1})}$ $=\mathrm{k}_{N}T_{11}-\mathrm{k}_{1}T_{22}-\mathrm{i}\left(T_{21}+\mathrm{k}_{1}\mathrm{k}_{N}T_{12}\right)\,,$ (27d)
Using similar arguments in terms of impulse response as we did with $\boldsymbol{T}$, we can think of $S_{11}$ as the outgoing response caused by an incoming amplitude unit impulse, while $S_{22}$ has the same meaning if the unit cell was traversed in the opposite direction. Thus, $S_{11}$ and $S_{22}$ must be equal (as the travel times do not change) except for the impedance contrasts being reversed to account for the opposite propagation direction. A similar interpreation goes for $S_{21}$ and $S_{12}$. The fact that all the entries $\boldsymbol{S}$ can be written in terms of those of $\boldsymbol{T}$ means that the former also admits a “path decomposition”, namely:
$S_{11}$ $=\sum_{j=1}^{2^{N-1}}\mathcal{S}_{j}e^{-\mathrm{i}\omega(\tau_{j}-t_{N}+t_{1})}\,,$ (28a)
$S_{22}$ $=\sum_{j=1}^{2^{N-1}}\mathcal{S}_{j}e^{\mathrm{i}\omega(\tau_{j}-t_{N}+t_{1})}\,,$ (28b)
$S_{12}$ $=\sum_{j=1}^{2^{N-1}}\mathcal{R}_{j}e^{-\mathrm{i}\omega(\tau_{j}-t_{N}-t_{1})}\,,$ (28c)
$S_{21}$ $=\sum_{j=1}^{2^{N-1}}\mathcal{R}_{j}e^{\mathrm{i}\omega(\tau_{j}-t_{N}-t_{1})}\,,$ (28d)
where the amplitudes $\mathcal{S}_{j}$ and $\mathcal{R}_{j}$ depend on the impedance of the interfaces and the wavenumber ratio between the first and the last layer:
$\mathcal{S}_{j}={1\over 2}\left(\mathcal{T}_{j}+{\mathrm{k}_{1}\over\mathrm{k}_{N}}\mathcal{T}_{j}^{\prime}\right)\,,\quad\mathcal{R}_{j}={1\over 2}\left(\mathcal{T}_{j}-{\mathrm{k}_{1}\over\mathrm{k}_{N}}\mathcal{T}_{j}^{\prime}\right)\,.$ (29)
Let us remark that in periodic media, $\mathfrak{Re}(S_{11})=\mathfrak{Re}(S_{22})=\eta$, the half-trace function that governs the dispersion relation in periodic 1D crystals. This is possible in periodic systems because one can always shift the unit cell to attain $\mathrm{k}_{1}=\mathrm{k}_{N}$.
4 Conclusions
We have outlined a combinatorial “path-by-path” procedure to directly construct all the entries of the cumulative transfer matrix $\boldsymbol{T}$ associated with an arbitrary number of layers $N$, and then used these to obtain the corresponding scattering matrix $\boldsymbol{S}$. Their physical meaning was previously concealed behind convoluted mathematical expressions that obscured an intuitive interpretation in terms of the various ways to traverse the layered system and the accompanying changes in amplitude and phase. The path decomposition (formerly termed “harmonic” decomposition) endows transfer matrices with a clear physical significance, so they do not have to be regarded just as a convenient computational tool.
Acknowledgements. The support of the Swiss National Science Foundation via Ambizione Grant 216341 “Data-Driven Computational Friction” is gratefully acknowledged.